OpenStudy (anonymous):

What is the sum of the first 55 positive multiples of 3? Please show your work.

OpenStudy (anonymous):

\[3+6+9+\cdots+162+165\] \[=3(1)+3(2)+ \cdots + 3(54)+3(55)\] \[=3(1+2+3+ \cdots +54+55) = 3 \cdot \frac{55(56)}{2}=4620\]

OpenStudy (anonymous):

Nice! I should've thought of that one. Thanks